# Detecting Multimedia Generated by Large AI Models: A Survey

### v1

@@ 5 Discussion
In this section, we discuss the main challenges and limitations encountered in the era of detection based on our comprehensive review of existing works. Then we propose potential directions for future research aimed at developing more efficient and practically effective detection systems.
@@ 5.1 Challenges
$\clubsuit$ Limited Reliability of Detection Models. A reliable detector should retain promising generalizability to new emerging models or data sources of detection, good interpretability in building trust in prediction results, and robust resistance against different real-life application conditions [251]. Therefore, we pinpoint the reliability challenges in detection models from three aspects: generalizability, interpretability, and robustness.
  • Generalizability. We can see from Fig. 1, that LAIMs and their generated multimedia are getting actively updated every three months. However, many current detectors are tailored to learn specific features from their training data, leading to overfitting and a lack of flexibility in detecting unseen data. For example, [44, 182, 183, 48, 184, 163] all highlight such generalization issues in detectors, especially in  [44, 163], both of them aim to simulate real-world scenarios. [44] demonstrates that commonly-used detectors have difficulty in detecting text in the wild. Similarly,  [163] shows that the detector’s performance drops substantially when the testing images come from generators that have major architectural changes from those that generate training data. All of those findings reveal the difficulty of reliable generalization to new emerging models or data sources of detection.
  • Interpretability. As multimedia generated by LAIMs becomes increasingly realistic and indistinguishable from human-created content, there is a growing reliance on methods that are easily understandable and tractable [252] instead of those with high accuracy but derived based on data that are hard to explain. Efforts like [190, 38] employ SHAP [221] to analyze local explanations of model’s decisions. However, their explanations are data-dependent.  [162] employs prototype-based explanation method, which can be susceptible to spurious correlations (confounding bias) [253]. The above methods still do not accurately reflect a detector’s decision-making process. Besides the model architecture, training data is another challenging aspect to consider, particularly in determining which specific features or subsets of the vast training data contribute to a detector’s predictions.
  • Robustness. Potential attacks are an important factor in the continued unreliability of current LAIM-generated multimedia detectors. These detectors are tasked with not only identifying content produced by the latest LAIMs but also demonstrating resilience against sophisticated attacks created by the most advanced generative models. [254, 255] illustrate this vulnerability, where they use diffusion-based models to craft attacks that successfully exploit weaknesses in existing detectors. Furthermore, [110] has shown that the performance of detectors is significantly influenced by image transformations. This finding emphasizes the challenges in ensuring the robustness of detectors, especially in real-world scenarios where images often undergo compression or blurring. Additionally, LAIM-generated text detection faces paraphrase attacks and prompt attacks performed by LAIMs, besides adversarial attacks [18, 19]. In a nutshell, LAIM-generated attacks and real-life detection scenarios bring a greater challenge to the robustness of detectors ever than before.
$\clubsuit$ Constrained Size, Accessibility, Range, and Variety of Datasets. As illustrated in Table I, there is a significant gap in the availability of large-scale datasets for video, audio, and multimodal detection. While there are a few million-scale image detection datasets, many have limited sample sizes (some are even less than ten thousand). Additionally, there is a noticeable lack of diversity in datasets. Current public multimodal datasets focus primarily on image and text modalities, overlooking valuable information that could be derived from other modalities like video and audio. Furthermore, the majority of text detection datasets are predominantly in English, and are limited to specific domains without encompassing a broad range of topics. The available fake audio dataset, LibriSeVoc [103], contains only two diffusion-based vocoders. This lack of diversity and narrow range lead to significant out-of-distribution challenges, causing detectors to underperform markedly when faced with data in real-world settings.
$\clubsuit$ Lack Evaluation Benchmark. The field of LAIM-generated multimedia detection faces the challenge of lacking a standardized benchmark, leading to inconsistent data processing, evaluation strategies, and metrics. This results in unfair comparisons and potentially misleading outcomes. For instance, Kamat et al. [152] note that detectors often tested on datasets similar to their training data, while Cocchi et al. [110] demonstrates that different data augmentation has notable impacts on the performance of detectors. Furthermore, the capability of LAIMs to partially manipulate multimedia content, such as through image inpainting, polishing human-written text, and synthesizing specific video frames, adds to the evaluation complexity. Therefore, how best to evaluate such intricately manipulated content with a unified benchmark remains challenging.
@@ 5.2 Future Directions
$\clubsuit$ Building Foundation Models. The term “foundation models” was first introduced by Bommasani et al. [256], and they defined foundation model as “the base models trained on broad data that can be adapted to a wide range of downstream tasks”, (e.g., CLIP [138], DALL-E 3 [120], ViT [139], and GPT family) . Multimedia generation has recently witnessed remarkable progress fueled by foundation models [4]. In LAIM-generated text detection, foundation models are widely used. Most recently, in fake image detection [150, 151, 149, 143, 162, 164], the paradigm is also shifting towards foundation models, because it allows rapid model development and provides better performance in real-world scenarios with their extensive pre-training on diverse datasets.
However, the application of foundation models in detection tasks remains quite limited. For instance, the above efforts in synthesized image detection are primarily limited to CLIP [138] and ViT [139]. This limitation is even more pronounced in areas like audio, video, and multimodal forgery detection, where there is a notable scarcity of research utilizing or developing foundation models. This gap indicates a significant opportunity for advancement in the field, particularly in integrating and exploring existing foundation models to bolster the detection capabilities of a wide range of digital forgeries. More importantly, there is a pressing need to develop multimodal detection-oriented foundation models. Such models would offer more robust and versatile detection systems compared to task-specific models, enabling researchers to rapidly adapt them for various downstream detection tasks, streamlining the detection process, and enhancing its efficacy.
$\clubsuit$ Harnessing Generative Discrepancies. Recent studies reveal clear differences between LAIM-generated multimedia and human-created content. LAIM-generated text, as shown by [50, 257, 258], often exhibits longer, more deterministic, and complex syntax compared to human writing. In image generation, current DMs face challenges in producing high-frequency information[88, 171], rich textures [94], repetitive patterns [177], and adherence to physical principles like lighting [179]. Effective detection strategies might involve a mix of these indicators, with potential in ensemble learning  [259], foundation models  [256], multi-modal/task learning, and cross-modal feature embedding.
@@
$\clubsuit$ Ensuring Reliability of Detectors. As reliability challenges in detection methods are discussed above, we urge forthcoming detection studies to prioritize detector reliability. The development of a reliable detector should aim for a unified framework that ensures generalizability, robustness, and interpretability concurrently. This could be a promising direction for detectors to combat LAIM-generated multimedia in real-world scenarios and increase the trust of the social community in detection results. Training on a diverse and large-scale dataset, building robust detectors with attacks, incorporating explainable AI (XAI) techniques [252], considering strategies like online learning [163] or transfer learning with new data to keep the model updated would contribute to building a reliable detector.
$\clubsuit$ Establishing Fair Detectors. It has been revealed that current detectors, though with high detection accuracy, result in unfair performance disparities among demographic groups, such as race and gender [260, 261]. This can lead to particular groups facing unfair targeting or exclusion from detection, potentially allowing misclassified LAIM-generated multimedia to manipulate public opinion and undermine trust in the model. Despite its criticality, the issue of fairness in LAIM-generated multimedia detection, even in detecting deepfake generated by traditional methods, has not received adequate attention from the research community. We advocate for a proliferation of studies that delve into addressing the bias problem in detectors. This involves not only developing a fair detector, but also ensuring that this fairness is generalizable across various scenarios without showing bias against certain groups.
$\clubsuit$ Accelerating the Development of Multimodal Detectors and Datasets. Existing research on detecting LAIM-generated multimedia mainly focuses on single data modality. However, in real-world scenarios, it is very hard to know which data modality is generated. For example, generated videos may have both visual and audio components generated, or only one modality may be created. Hence, the creation of multimodal detectors is crucial. To effectively develop these models, assembling extensive multimodal datasets for their training is a fundamental requirement. This can be explored in the future.
$\clubsuit$ Towards Interpretable User-friendly Detection Tools. Our analysis of popular online detection tools discussed in Section 4, reveals a common shortcoming: a lack of interpretability. However, humans generally prefer tools that are directly interpretable, tractable, and trustworthy [252]. Developing interpretable detection tools can gain more social attention and popularity. Therefore, enhancing interpretability is not just a technical upgrade but a significant step in making these tools more accessible to broader users, especially beneficial for underrepresented groups such as non-technical persons and elderly adults.

### v4

@@ 5 Discussion
In this section, we summarize shared detection strategies and key insights across modalities. We further discuss both common and modality-specific challenges, with particular emphasis on real-world environments such as social media. Guided by our analysis, we outline promising research directions in a modality-specific manner that correspond to addressing these challenges.
5.1 Shared Techniques and Cross-Modal Insights
Despite differences in data type, detection of AI-generated content often relies on a similar toolkit of strategies.
$\clubsuit$ Statistical Anomaly Detection. AI-generated content often exhibits characteristics that deviate from those of human-produced data. For instance, LAIM-generated text tends to be more objective and formal, less prone to biased or harmful language, and generally longer and more detailed than human-authored text [295, 402], and artifacts appear in co-occurrence patterns [325]. Similarly, LAIM-generated images frequently contain frequency domain artifacts (e.g., periodic noise from upsampling) that diverge from the statistical regularities of natural photographs [102, 97]. In the audio domain, researchers similarly exploit spectrogram inconsistencies and vocoder-specific artifacts to distinguish synthetic speech from human speech [132, 240]. A recent work [271] draws direct inspiration from LLM-based text detectors like DetectGPT [323], which evaluate the log-likelihood or "surprise" of a token sequence under the model’s learned distribution. Similarly, their zero-shot image detector (ZED) treats an image as "statistically surprising" if it deviates from the distribution modeled by a lossless encoder trained solely on real images. This analogy demonstrates how distributional surprise can serve as a modality-agnostic principle for synthetic content detection, manifesting as log-perplexity in text, and as coding cost gaps in images. Across modalities, detection systems exploit these distributional anomalies by identifying statistical irregularities or generative artifacts.
$\clubsuit$ On the Use of Model Fingerprints and Manifold Traces. Recent advances across modalities reveal that generative models often exhibit statistically or geometrically structured biases in their outputs, which can be exploited through perturbation or reconstruction-based analysis. In text detection, a growing body of work has used log-probability curvature, the second-order change in log-likelihood under small input perturbations, as a robust indicator of LLM-generated text. For example, DetectGPT [323] observes that AI-generated text tends to lie on flatter regions of the model’s learned probability manifold and proposes a curvature-based detection criterion. Later works improve efficiency using Bayesian surrogate models [315], conditional sampling [316], and sentence-level token probability aggregation [318]. Analogously, diffusion fingerprinting techniques in image detection leverage the structured nature of generative image manifolds. Methods such as DIRE [93] and SeDID [278] measure how well a pre-trained diffusion model can reconstruct an input image. The hypothesis is that generated images sampled directly from the model’s distribution exhibit lower reconstruction error than real images. This mirrors the principle of log-probability curvature: real data tends to fall off the generative manifold, leading to higher deviation under model-guided reconstruction. Brokman et al. [281] formalize this intuition by defining curvature and gradient-based geometric criteria on the score function manifold of a diffusion model. These quantities, computed over perturbed versions of the image, serve as zero-shot indicators of generation. The shared insight across modalities is that the local geometry of the model’s generative manifold, whether measured via reconstruction error or curvature of probability estimates, encodes telltale signs of synthetic content. These techniques, rooted in manifold analysis, suggest promising future directions for cross-modal detection frameworks that generalize by probing the generative "fit" of a given sample rather than relying on domain-specific artifacts alone.
$\clubsuit$ Detect Watermark and Embedded Signals. A widely adopted detection oriented strategy across modalities is to detect imperceptible signals, known as watermarks, that are embedded into AI-generated content at generation time. In text, this typically involves subtly biasing token selection toward a secret subset of vocabulary based on a hidden key, producing statistical patterns detectable through lightweight hypothesis tests without compromising fluency [301, 334, 331]. In images, watermarking advances embed signatures directly into the latent representations of diffusion models, allowing detection even after transformations like cropping or compression [286, 403]. Detecting watermarks provides a modality-agnostic and efficient means of verifying AI-generated content. Beyond binary classification, watermark-based methods also support model attribution, enabling the identification of the specific generative model responsible.
$\clubsuit$ Disentangled Representation Learning for Generalization. One technique transferring across modalities is disentangled representation learning aimed at isolating AI-generated artifacts from the underlying content. In image deepfake detection, researchers found that naive classifiers often overfit onto content features rather than the forgery artifacts, hurting generalization to new data [404, 405]. To address this, Liang et al. [405] proposed a framework that disentangles image features into content and artifacts, training detectors solely on the latter. A similar approach has since been applied to LAIM-generated audio. Ren et al. [240] propose a disentanglement framework for synthetic speech, using multitask and contrastive learning to isolate vocoder-agnostic artifacts. This improves generalization across voice vocoders, with only vocoder-agnostic artifact features used for final classification. Extending this idea to text, Bhattacharjee et al. [303] propose ConDA, a contrastive domain adaptation framework that separates generator specific patterns from task relevant features. By training on labeled and unlabeled data from different models, it learns domain-invariant representations that generalize to unseen text generators.
$\clubsuit$ Semantic Context and Reasoning. The use of high-level reasoning was traditionally the realm of text (logic and consistency in narrative). Now, image and video detection are incorporating semantic reasoning (with the help of MLLMs) [232, 228, 227, 110, 120], effectively importing NLP-style consistency checks into visual analysis. Conversely, text detection is starting to use “world knowledge” (via models) to ask if a stated fact is real or a hallucination (which could hint at AI generation) [406]. This blending of contextual, semantic analysis means the line between modalities is blurring: detecting a fake might involve simultaneously reading text, analyzing image pixels, and using background knowledge from a knowledge base.
@@ 5.2 Challenges
5.2.1 Common Challenges Across Modalities
$\clubsuit$ Generalization to New Generators. A persistent challenge is the limited generalizability of detectors to content from novel large AI models [63, 114, 136]. Generative models evolve rapidly, but many detectors overfit to artifacts of the training data and struggle to detect content generated by new generators.
$\clubsuit$ Robustness Against Adversarial Attacks and Post-Processing. Detection methods, regardless of the modality they target, currently struggle with adversarial manipulations, paraphrasing, and common post-processing operations, such as compression, resizing, which often obscure or erase the subtle artifacts that detectors rely on. Detected text can be paraphrased or lightly edited to evade classifiers [30, 293], and audio deepfakes can be passed through filters to mask telltale traces [136]. Minor perturbations, a slight rotation or noise addition to an image can fool many detectors [249, 246].
$\clubsuit$ Open-Domain and Social Media Complexity. The detection of LAIM-generated content across modalities on open-domain platforms, such as social networks, presents significant challenges due to the informal, noisy, and heterogeneous nature of social media content [155, 110, 112]. For text, posts are brief, laden with slang, emojis, or hashtags, diverging from the formal prose that detectors are usually trained on. This domain mismatch means a model effective on well-written text may fail on a jargon-filled tweet. Likewise, user-generated images and videos often contain filters, edits, or montages; one post might be entirely LAIM-generated while another only has an LAIM-edited segment. Consequently, the complexity and variability of open-domain social media content pose obstacles for detectors trained on idealized, tightly controlled laboratory data [130], necessitating more robust and adaptable detection frameworks.
$\clubsuit$ Lack of Interpretability. Across all modalities, most detection systems operate as black boxes, providing binary real/fake outputs without human understandable explanations. This lack of transparency makes it difficult for users, moderators, and policymakers to trust or validate detection outcomes, especially in high stakes contexts like content moderation, journalism, or forensic analysis. While some recent efforts (e.g., IPAD [407] for text and SIDA [110] for image) have begun incorporating explainable outputs, such as generating inverse prompts to reveal likely AI intent or highlighting tampered regions in images, these methods still fall short of accurately reflecting the detector’s underlying decision making process or providing comprehensive, human interpretable justifications for their judgments.
5.2.2 Modality-Specific Challenges
$\clubsuit$ Text-Specific Challenges. AI-generated text detection faces unique challenges in terms of generalization, robustness, and interpretability. A key issue in generalization is the limited cross-lingual and domain adaptability of current detectors. Many detectors are trained primarily on English datasets, making them less effective on non-English texts [43]. Similarly, detectors trained on domain-specific datasets, such as academic articles, may struggle when applied to other domains like social media posts, news articles [34]. From a robustness standpoint, paraphrasing attacks remain a major vulnerability [62]. Light edits, such as synonym replacement, sentence restructuring, or rewording via another LLM, can obscure the statistical patterns or watermark signatures that detectors rely on, allowing LAIM-generated content to bypass detection even when its meaning remains unchanged.
Attribution poses additional difficulties. Current attribution models often suffer from domain dependence, showing sharp performance drops on unseen styles or topics [307, 168]. Furthermore, fine-tuning and sampling variability introduce ambiguity: the same underlying model can produce stylistically diverse outputs based on fine-tuning or sampling parameters, making attribution even harder [313] The growing number of LLM variants, both open and proprietary, also creates scalability challenges[27, 314], and representation overlap across models reduces discriminability[312].
Moreover, interpretability is notably lacking in most text detection systems. Many models offer binary predictions without articulating the reasoning behind their decisions, which limits their transparency and trustworthiness. Although recent approaches like IPAD [407] attempt to improve interpretability by inferring a plausible inverse prompt that explains the text’s generative origin, such methods are still in early stages and not widely adopted.
Social Media. Models trained on formal English text, such as news articles, often struggle with informal, multilingual, or domain-shifted content like social media posts or low-resource languages, revealing poor performance across diverse linguistic and contextual domains [34, 63]. The informal, brief, and multilingual nature of social media posts, often laden with slang, emojis, hashtags, or internet jargon, exacerbates this domain shift, rendering detectors trained on formal or long form data less effective on platforms like X or TikTok [66]. Recent benchmarks like MultiSocial [66] demonstrate that fine-tuning detectors on social media data can improve performance, but this success is often platform or language-specific, requiring retraining to maintain effectiveness across different platforms or languages.
$\clubsuit$ Image-Specific Challenges. Detecting LAIM-generated images remains challenging due to the increasing visual fidelity of modern generative models and the fragility of detectors under real-world conditions. In addition to the common challenges discussed above, the image modality faces specific obstacles related to robustness against post-processing and partial manipulations, localization and attribution, limited interpretability of detection outputs, and growing concerns about demographic fairness.
Robustness is a critical issue, benchmarks demonstrate that even mild JPEG compression or resolution reduction significantly degrades detection accuracy [116, 114]. Social media platforms exacerbate this challenge by routinely applying such transformations to uploaded images, necessitating highly robust detectors capable of handling diverse, processed content.
Beyond binary detection, localization itself presents additional challenges. Pinpointing partially manipulated regions is harder than global detection, as edits are often small, well-blended, and vary by forgery type [92]. Fine-grained ground truth is also labor intensive to obtain [254]. Moreover, artifacts tend to cluster near object boundaries or semantic inconsistencies [255], motivating multi-view or object-level modeling [256]. Attribution, which identifies the generative model behind a fake image, faces several challenges. It is highly sensitive to domain shifts, minor post-processing like compression or resizing can degrade performance [236, 237]. Attribution models also struggle to generalize to unseen or updated generators, as the fingerprints they rely on often change [270]. Unlike binary detection, attribution requires multi-class classification, increasing model complexity and the risk of confusion between similar generators.
Interpretability also remains limited. Most vision-based detectors output binary labels or heatmaps that highlight suspicious regions but do not provide clear, human-readable explanations. While recent frameworks like SIDA have introduced explanation-aware detection by combining tampered region localization with textual justifications, such methods are not yet standard and often struggle to align model reasoning with human perceptual cues. Lastly, fairness in image detection is an emerging concern [116], as detectors may exhibit bias across demographic groups. For example, forensic artifacts can manifest differently depending on skin tone, facial features, or lighting conditions, potentially leading to disparate false positive rates across racial or gender groups.
Social Media. Common social media operations, such as compression, resizing, or filtering, can erase subtle artifacts or noise patterns that detectors rely on. As a result, detectors trained on clean, high-quality datasets often experience severe performance degradation when applied to social media content [110, 112]. Another challenge is that LAIM-generated images shared on social media are often open-domain, covering a wide range of subjects beyond specific categories like faces or objects. Detectors trained on domain-specific datasets, such as face-focused datasets [116], may fail to generalize to diverse content types in the open-world setting.
$\clubsuit$ Video-Specific Challenges. In addition to inheriting the challenges faced by image-based detection, video introduces added complexity due to its temporal dimension. Convincing LAIM-generated deepfake videos often exhibit smooth motion, consistent lighting, and accurate lip-sync across frames, leaving only subtle temporal artifacts, such as unnatural blinking, jitter, or slight misalignments between facial expressions and speech, for detectors to identify. Additionally, the development of robust and generalizable video detection methods is hindered by the lack of large-scale, diverse datasets featuring LAIM-generated videos. Furthermore, the absence of standardized video deepfake benchmarks makes it difficult to compare methods fairly or assess their generalization to real-world deployment scenarios. Lastly, video is inherently multimodal, typically accompanied by audio that may itself be synthesized. This requires detection systems to go beyond visual cues and incorporate cross-modal consistency analysis, such as verifying alignment between speech and lip movement or between speaker identity and voice.
Social Media. Social media platforms further complicate detection. Same as images, compression, resolution reduction, filtering, and watermarking, common on platforms like TikTok and X, can obscure the subtle forensic artifacts that detectors rely on. Moreover, the fast-paced, high-volume nature of social media demands near instantaneous analysis to flag LAIM-generated videos before they spread widely, yet processing high-resolution, temporally complex, and multimodal content in real-time remains computationally demanding. Current detection systems, often trained on curated datasets, struggle to achieve the speed and accuracy needed for real-time deployment.
$\clubsuit$ Audio-Specific Challenges. Detecting LAIM-generated audio, such as synthetic speech or voice clones, is particularly difficult due to the subtlety of generative artifacts and the variability of real-world acoustic conditions. Forensic cues, such as unnatural prosody, spectrogram anomalies, or phase irregularities, are often very subtle and difficult to detect consistently. The broad range of recording qualities, from clean studio recordings to noisy mobile captures, exacerbates the generalization problem, as models trained on curated datasets often fail in uncontrolled settings. This challenge is compounded by the lack of large-scale, diverse LAIM-generated audio datasets and standardized benchmarks, which hinders both model development and fair evaluation across methods.
Social Media. Social media platforms add additional challenges for LAIM-generated audio detection. Audio shared online is often compressed into low-bitrate formats, mixed with background noise, or embedded within videos, further masking forensic traces. Compression, platform specific filtering, and varying playback conditions can distort or erase subtle synthetic artifacts, making detection much harder. As with video, the fast dissemination speed on social media requires detectors to operate in real time, balancing accuracy with computational efficiency under noisy, degraded conditions.
$\clubsuit$ Multimodal-Specific Challenges.LAIM-generated multimodal content, where text, images, audio, and video are combined, presents distinct detection difficulties beyond those seen in unimodal scenarios. One challenge lies in identifying cross-modal inconsistencies, such as mismatches between audio and lip movements or between image content and accompanying text. Single modality detectors often fail to capture such discrepancies. Further complicating detection is the presence of partially manipulated media, where only one modality (e.g., audio in a real video) is synthetic, requiring systems to localize which component is fake. Moreover, there is a critical lack of large-scale LAIM-generated multimodal datasets and standardized benchmarks that cover richly entangled combinations of modalities. Recent evaluations, such as those in the LOKI benchmark [154], show that no universal detection model currently performs reliably across modalities: large multimodal models, while promising in reasoning and explainability, still exhibit modality imbalance and perform poorly on complex tasks involving audio or video. These challenges highlight the nascency of multimodal detection research and the urgent need for scalable, explainable, and cross-modality aware detection systems.
Social Media. On social media platforms, multimodal content is especially prevalent, often subject to compression, stylization, or casual editing. These platform specific transformations obscure cross-modal cues and make it easier to conceal manipulations across one or more modalities. For example, AI-generated deception can subtly modify a single modality while preserving overall narrative coherence [146]. Furthermore, the need to analyze multiple data streams (e.g., video frames, audio tracks, and text captions) imposes substantial computational overhead, which is non-trivial at the scale of social media.
5.2.3 Challenges of Detection Tools
A primary challenge lies in the lack of interpretability and user trust. Many current detection tools operate as black boxes, offering binary outputs ( or confidence scores without any contextual explanation or rationale. This lack of interpretability can undermine user trust, particularly for non-technical users such as journalists, educators, or fact-checkers, who require transparency to make informed decisions.
A second limitation is the lack of transparency. As shown in Table II, the majority of high performance tools are closed-source commercial platforms. These systems typically do not disclose details about their model architectures, training datasets, or potential biases, which hinders scientific reproducibility and independent evaluation.
Lastly, current tools often fall short in supporting real-time detection. Most existing tools are only accessible through web applications, which may not align well with user needs for fast and real-time detection. For example, users may need to analyze audio or video content on social media platforms in real time, but most current tools often require uploading files manually through dedicated portals.
@@ 5.3 Future Directions
5.3.1 Common Future Directions Across Modalities
@@
$\clubsuit$ Enhance Generalizability, Robustness, and Interpretability. Future detection systems must be more adaptable to evolving generative models, robust against real-world distortions, and transparent in their decision-making. Training on a diverse and large-scale dataset, building robust detectors with attacks, incorporating explainable AI (XAI) techniques [408], considering strategies like online learning [259] or transfer learning with new data to keep the model updated would contribute to building such a reliable detector.
$\clubsuit$ Social Media Adaptation. As LAIM-generated content proliferates on platforms like X, TikTok, and Reddit, future detection systems must adapt to the unique constraints of social media environments. These platforms introduce compression, format conversions, and short form or informal content that degrade detection performance [66, 241, jiang2024localizing]. To enhance robustness, future work should incorporate platform specific data and leverage domain adaptation or robustness oriented training strategies [250, 409]. Additionally, integrating rapid provenance verification mechanisms, such as watermark detection [286, 410] or cryptographic signatures like C2PA [411], can enable scalable real-time screening prior to deep forensic analysis. Several studies also highlight the need for social-media aware benchmarks to reflect realistic deployment conditions and avoid overfitting to clean or curated datasets [67, 343, 224]. Across modalities, social media adapted detection remains a crucial direction for practical, deployable solutions.
5.3.2 Modality-Specific Future Directions
$\clubsuit$ Text–Future Directions. Advances in LAIM text detection will center on making models generalizable across different languages. One approach is to train multilingual detectors so that detection is not limited to English texts. Researchers are also investigating fine-grained stylometry and semantic consistency checks, for instance, does the text follow unnatural repetition or omit context in a way that humans typically would not? Addressing the text domain’s challenges requires embracing its diversity: models must become fluent in the spectrum of human and AI writing styles to flag machine-generated content reliably in any context.
Social Media. Future detectors should also learn from social media data, mastering the idioms, shorthand, and irregular grammar that characterize those platforms. Incorporating training corpora from social media and building models that recognize human-like variability versus AI consistency will be critical for improving robustness. Given the casual and often fragmented nature of social media language, detectors must balance sensitivity to AI artifacts with tolerance for authentic human irregularities.
$\clubsuit$ Image – Future Directions. As LAIM-generated images become increasingly photorealistic, future detection systems should integrate diverse forensic cues into a unified framework. Despite recent advances, current diffusion models still exhibit limitations in generating high frequency details [97, 272], rich textures [103], repetitive patterns [282], and physically consistent lighting [284]. Effective detection strategies may leverage a combination of these shortcomings by incorporating multiple complementary signals. For example, frequency domain analysis can detect hidden spectral anomalies or periodic artifacts subtle generative “fingerprints” that persist even in natural looking images. Physics-based cues, such as inconsistencies in lighting, shadows, and reflections, can expose violations of real-world constraints that are difficult for generative models to replicate. Physiological cues, such as unnatural facial geometry or distorted proportions, further contribute to both robustness and interpretability. These heterogeneous cues can be effectively utilized through techniques such as ensemble learning[412], multitask learning, or by employing foundation models[413] trained across diverse visual forensic tasks.
Social Media. Rapid provenance verification, checking for cryptographic certificates of origin or embedded authenticity signals, will become increasingly important for detecting social media processed images. Initiatives like C2PA [411] aim to attach secure provenance metadata to images at creation time, enabling detectors to validate content before resorting to computationally intensive forensic analysis. Similarly, watermarking strategies [286, 410] are being developed to embed resilient signals that survive typical social media transformations. Future detection pipelines may thus adopt a two-tier approach: a fast check for provenance indicators (metadata or watermarks), followed by deeper forensic investigation if no authenticity signals are found. This strategy could enable scalable monitoring and early flagging of LAIM-generated images in the high-volume, fast-paced environment of social media.
$\clubsuit$ Video – Future Directions. Future research should explore finer-grained modeling of temporal dynamics across frames. This could include the design of lightweight yet expressive spatiotemporal architectures that integrate intra and inter-frame attention mechanisms to capture inconsistencies in motion continuity. Such modeling is particularly important given that even advanced generative models like Sora often exhibit temporal artifacts such as jittery motion, abrupt transitions, and object flicker [414, 415]-failures that cannot be reliably detected by frame-level analysis alone. Real-time detection remains a bottleneck due to the high computational cost of analyzing full resolution video streams with temporal context. To meet this demand, future work should prioritize model compression techniques (e.g., quantization, pruning, knowledge distillation) and develop efficient early-exit mechanisms or key-frame selection strategies to reduce latency. Architectures optimized for edge inference can facilitate on-device deployment on mobile or streaming platforms. Finally, the absence of large-scale, diverse, and standardized datasets of LAIM-generated videos hinders both generalization and benchmarking. Future initiatives should aim to create dynamic, open-source datasets that encompass a range of generative models, scene types, and manipulations. Alongside this, establishing unified evaluation benchmarks and protocols would enable consistent assessment of detection performance across varying video conditions, promoting methodological comparability and real-world readiness.
Social Media. Compression and resolution degradation applied by social platforms can distort forensic traces. One promising direction is the development of preprocessing pipelines that learn to restore or enhance compression-corrupted artifacts relevant to forgery detection [416, 417]. Additionally, adversarial training and domain adaptation strategies [409, 418] could improve generalization across post-processed and platform specific video formats. Integrating rapid provenance verification, such as detecting watermark signatures or metadata if available, before deep forensic analysis could further help scale detection systems.
$\clubsuit$ Audio – Future Directions. One of the biggest challenges in detecting LAIM-generated audio is the lack of large-scale, diverse datasets and standardized benchmarks, which significantly hinders the development and fair evaluation of audio detection systems. To address this, future research should expand the scope of training data to include a wide variety of synthetic audio samples spanning multiple languages, speaker identities, generation techniques, and real-world distortions such as background noise or channel degradation. Exposure to such diversity will improve model generalization and robustness in uncontrolled environments. Beyond dataset expansion, future efforts should also focus on more robust forensic features. Rather than relying solely on spectral or vocoder artifacts, which are often erased by compression or playback, detectors may increasingly target features that are harder for AI to replicate, such as breathing patterns [419], micro-pauses [420], and subtle pitch or timing irregularities [421] that emerge from human cognitive and physiological processes.
Social Media. Audio deepfake detection on social media faces unique challenges, as audio is often shared alongside video and subjected to platform specific compression and noise. A promising direction is the use of speaker verification and authentication techniques that leverage speaker embeddings to capture individual-specific vocal traits [422, 423]. By incorporating speaker verification frameworks, detectors can perform personalized authenticity checks, comparing a user’s voice against reference samples to detect mismatches indicative of manipulation. On social media, identity-aware methods that jointly analyze audio and visual cues, such as fusing speaker recognition with facial analysis, can further enhance detection robustness, particularly when cross-modal inconsistencies between voice and appearance are present.
$\clubsuit$ Multimodal–Future Directions. Future multimodal detection systems should jointly analyze text, image, audio, and video to uncover cross-modal inconsistencies and detect partial modality manipulations. This calls for unified or modular architectures capable of both localizing manipulated modalities and reasoning across them. Developing such systems requires large-scale, LAIM-generated multimodal datasets that reflect realistic, partially edited content. Recent benchmarks such as LOKI [154] and MMFakeBench [155] reveal that current large multimodal models still struggle, particularly on audio and video inputs, underscoring the need for modality-balanced training and fine-grained supervision. Advances in MLLMs (multimodal large language models), such as LLaMA 4 [424], GPT-4o [214], and Gemini 2.5 Flash [425], combined with domain-specific forensic knowledge, offer promising foundations for building robust multimodal detection systems.
Social Media. To enable reliable verification in complex, real-world scenarios such as social media, future multimodal detectors must be scalable and resilient to noisy, user-generated content, including informal, stylized, or culturally nuanced formats like memes and TikTok videos. Achieving this may require specialized pre-processing pipelines and platform specific training strategies. Rapid provenance verification and modality-specific robustness enhancements will also be important for scalable deployment.
5.3.3 Future Directions for Detection Tools
Our analysis of popular online detection tools discussed in Section 4.1, reveals a common shortcoming: a lack of interpretability. Developing interpretable detection tools can therefore gain more social attention and user adoption. To enhance interpretability, future systems could integrate real source verification, physical and physiological cues, and multimodal large language model reasoning to provide contextual explanations for why content is identified as synthetic. Moreover, human-in-the-loop detection platforms (e.g., involving expert review of challenging cases) that combine AI analysis with expert review can ensure more accountable deployment. To improve transparency and reproducibility, open-sourcing model architectures, datasets, and evaluation protocols, where ethically and legally feasible, should be prioritized. In addition, to better support real-time detection needs, future tools should be designed with lightweight, integrable formats such as browser extensions, enabling seamless deployment across social media, messaging apps, and content platforms. This is not only a technical upgrade but also makes these tools more accessible to broader user groups, particularly underrepresented populations such as non-technical individuals and elderly adults.